# Modeling the amplification of epidemic spread by misinformed populations

### v2

@@ Appendix 1 Mean-field SMIR model
We now explore the effect of the recovery rate, again in the scenario with no misinformation or homophily, by setting $\beta_{O}=0.3$ and varying the recovery period $\tau=1/\gamma$ between 1 and 20 days. Fig. S2 shows that when $\tau<4$, $R_{0}<1$ and the disease does not reach epidemic proportions. At this level, the epidemic takes a long time to reach its peak ($\approx$80 days). Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly. We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC14. With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
@@ Appendix 2 Robustness analyses
We also investigate the impact of the contact network density ($\bar{k}$) on infection dynamics, as illustrated in Fig. S4 for $5\leq\bar{k}\leq 25$. We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions30. As expected, Fig. S4(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission. But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases. This is because, at higher $\bar{k}$ values, the infected population is already substantial even in the low-misinformation ($\phi=20$) baseline. The combined effect of these two opposing trends, as shown in Fig. S4(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$. Fig. S4(b) also shows that, consistent with our primary findings, larger $\phi$ (misinformed resilience) results in a smaller infected population. In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts, resulting for example from social distancing or lockdowns, by decreasing the $p$ parameter.

### v3

@@ Acknowledgments
Author contributions
The initial concept for this research was developed by YA, SF, AF, FM. The final study design was developed collectively by all authors. Data collection was led by MRD as part of the CoVaxxy project 21. All mean-field modeling was conducted by MRD. FP developed the early code for the agent-based simulation, which was subsequently refined and expanded by MRD for publication. Analysis was conducted primarily by MRD, with help from FP, and guidance from YA, SF, AF, FM. The SMIR model was developed collaboratively by all authors. Visualizations were created by MRD, FP, and FM with input from all authors. The first draft was written by MRD, with revisions from all authors. FM oversaw the progression of the study.
Competing interests
Authors declare that they have no competing interests.
@@ Appendix 1 Mean-field SMIR model
We now explore the effect of the recovery rate, again in the scenario with no misinformation or homophily, by setting $\beta_{O}=0.3$ and varying the recovery period $\tau=1/\gamma$ between 1 and 20 days. Fig. S2 shows that when $\tau<4$, $R_{0}<1$ and the disease does not reach epidemic proportions. At this level, the epidemic takes a long time to reach its peak ($\approx$80 days). Increasing $\tau$ means that individuals remain infected longer, so the population gets infected faster and the peak infection is reached more rapidly. We set $\tau=5$ ($\gamma=0.2$) in the main text to align with quarantine recommendations from the CDC15. With this value and $\beta_{O}=0.3$, the dynamics are at epidemic levels ($R_{0}=1.5$).
@@ Appendix 2 Robustness analyses
We also investigate the impact of the contact network density ($\bar{k}$) on infection dynamics, as illustrated in Fig. S4 for $5\leq\bar{k}\leq 25$. We consider this range because $\bar{k}=25$ represents pre-pandemic daily social contacts while $\bar{k}=5$ represents COVID-19 lockdown conditions32. As expected, Fig. S4(a) demonstrates that higher $\bar{k}$ leads to increased infections through the population, since the higher contact density provides more opportunities for transmission. But while a larger percentage of the overall population is infected, the relative effect of misinformed individuals decreases. This is because, at higher $\bar{k}$ values, the infected population is already substantial even in the low-misinformation ($\phi=20$) baseline. The combined effect of these two opposing trends, as shown in Fig. S4(b), is that the additional percentage of infected individuals relative to the $\phi=20$ scenario reaches a maximum for some intermediate $\bar{k}$. Fig. S4(b) also shows that, consistent with our primary findings, larger $\phi$ (misinformed resilience) results in a smaller infected population. In our main analysis we focus on $\bar{k}=25$ and model an effective reduction of contacts, resulting for example from social distancing or lockdowns, by decreasing the $p$ parameter.